OpenStudy (anonymous):

HELP PLEASE Given the polynomial 6x3 + 4x2 - 6x - 4, what is the value of the coefficient 'k' in the factored form? 6x3 + 4x2 - 6x - 4 = 2(x + k)(x - k)(3x + 2) k= _____________

OpenStudy (campbell_st):

equate the constant terms \[-4 = 2 \times k \times -k \times 2\] solve for k

OpenStudy (anonymous):

can you just give me the answer? :(

OpenStudy (anonymous):

not so good at this

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

if k = 1 then -4 = -4 times 1^2